# Diffuse-CLoC: Guided Diffusion for Physics-based Character Look-ahead Control

### v2

@@
Abstract
We present Diffuse-CLoC, a guided diffusion framework for physics-based look-ahead control that enables intuitive, steerable, and physically realistic motion generation. While existing kinematics motion generation with diffusion models offer intuitive steering capabilities with inference-time conditioning, they often fail to produce physically viable motions. In contrast, recent diffusion-based control policies have shown promise in generating physically realizable motion sequences, but the lack of kinematics prediction limits their steerability. Diffuse-CLoC addresses these challenges through a key insight: modeling the joint distribution of states and actions within a single diffusion model makes action generation steerable by conditioning it on the predicted states. This approach allows us to leverage established conditioning techniques from kinematic motion generation while producing physically realistic motions. As a result, we achieve planning capabilities without the need for a high-level planner. Our method handles a diverse set of unseen long-horizon downstream tasks through a single pre-trained model, including static and dynamic obstacle avoidance, motion in-betweening, and task-space control. Experimental results show that our method significantly outperforms the traditional hierarchical framework of high-level motion diffusion and low-level tracking. We also show our results in the accompanied video: https://www.youtube.com/watch?v=kxf9dpv0dhw
I Introduction
@@
Recent advances in kinematic motion diffusion models have shown promise in motion synthesis through conditioning on partial states [1], text [2], and music [3]. Furthermore, the same kinematics motion diffusion model can be reused on unseen downstream tasks via inference-time conditioning such as classifier-guidance or inpainting [4, 5]. Yet, these approaches fall short of achieving physical realism. To improve physical consistency, a tracking policy can be used to follow the kinematics motions, but ensuring robustness to domain shifts introduced by the kinematics generation remains a challenge. On the other hand, prior works [6, 7, 8] have focused on predicting actions, ensuring physical realism without requiring extra components. However, unlike kinematic-based diffusion models, these approaches cannot adapt to unseen downstream tasks without retraining, as inference-time conditioning techniques are infeasible due to the lack of direct comparability between goals defined in state space and synthesized action trajectories.
@@
Our primary insight lies in modeling the joint distribution of states and actions within a single diffusion model, which enables action generation to be conditioned on the predicted states. Building on this insight, we develop a transformer-based diffusion architecture and an attention mechanism that includes non-causal attention for states and causal attention for actions. We further introduce a rolling scheme inspired by [9] to improve consistency and speed in auto-regressive policy execution. Ultimately, Diffuse-CLoC enables the application of established techniques from kinematic motion generation, such as classifier guidance, to perform multiple unseen downstream tasks with physics-based characters, including waypoint navigation, dynamic obstacle avoidance such as other characters, and dynamic maneuvers like jumping over obstacles and crawling underneath them, as well as inpainting for motion in-betweening task (Fig. LABEL:fig:teaser). Our primary contributions include:
@@ II Related Work
II-A Physics-Based Character Animation
@@
Early physics-based character animation focused on tracking single motion trajectories [10]. Subsequent work expanded this to capture motion distributions, such as AMP [11], though these approaches faced challenges in training stability and were limited to small datasets. As an alternative, latent skill embeddings were introduced, leveraging a VAE trained on large datasets [12, 13, 14, 15], which can also be paired with a high-level controller to complete downstream tasks [16, 15, 14].
@@
Recent motion tracking controllers [17, 18] have achieved notable success, leading to approaches that decompose the problem into two parts: a kinematic motion synthesis model and a universal tracking controller for the synthesized motion [19, 20]. However, the tracking controller’s performance heavily depends on the quality of the kinematic trajectory, which can include artifacts like floating bodies, foot sliding, and penetration.
@@
To address this challenge, RobotMDM [21] fine-tunes a motion diffusion model using the value function of an RL tracker to produce more trackable motions. Similarly, [22, 23] refine diffusion models directly with PPO-based reinforcement learning guided by an RL tracker’s reward. Other methods instead fine-tune the tracking controller to handle artifacts created by the diffusion model [20].
@@
Despite these advancements, trackers are often sensitive to out of distribution motions and require task-specific fine-tuning for applications like object interaction [20]. Furthermore, we demonstrate that even minor classifier guidance on kinematic trajectories can hurt the robustness of general tracking controllers.
II-B Diffusion in Physics-based Control
@@
Diffusion policies have been successfully applied in character animation [7], manipulation [6], and locomotion [8, 24] domains. However, its action-only generation makes inference-time conditioning impractical. In contrast, Diffuser [25] diffuses both states and actions, enabling classifier guidance through a reward formulation though exhibits limited robustness [26]. Decision Diffuser [27] improves robustness by similarly diffusing both states and actions, but discards actions during inference, relying instead on a trained inverse dynamics model. In contrast, our method enables both intuitive guidance directly in state space and high-quality conditional generation in action space.
II-C Steering Motion Synthesis
@@
Classifier-free guidance enhances sample quality and control by interpolating between an unconditional and a conditional model during the generation process, without requiring an explicit classifier [28]. In character animation, this method has been used to synthesize motions with text conditioning [7, 2]. However, the conditioning signals must be specified during training, and the model must be re-trained for any new conditioning signals.
@@
Classifier guidance incorporates conditional information after the training phase by utilizing the gradient of a pre-trained classifier [29] or cost function [30]. This technique guides the diffusion model’s generative process, steering outputs toward the desired class or minimizing the loss. In character animation, it allows a single pre-trained kinematics model to address multiple downstream tasks, such as obstacle avoidance and reaching [4, 25]. Despite its success, classifier guidance has not yet been explored for dynamic tasks in underactuated robotic systems, such as physics-based character animation.
@@
Inpainting involves generating or reconstructing missing regions of the original signal by conditioning the diffusion process on the surrounding known context. In character animation, this technique is useful for generating trajectories with partially known states, such as root trajectories [4] or a set of desired joint trajectories [3].
@@
Reinforcement learning has been employed as a high level planner to direct the diffusion process for downstream tasks [1]. When future trajectory predictions are absent, a high-level planner is essential to strategize [1]; however, our method can plan and foresee future events, allowing the accomplishment of downstream tasks without the need for a high-level planner.
II-D Inference Consistency and Speed
@@
A key challenge for generative planners is ensuring that the generated plans do not conflict during rollouts. Frequent replanning with multimodal trajectory distributions can lead to alternating behavior modes, negatively affecting task performance [31]. One approach is to execute a fixed action sequence before replanning [6], but this lacks real-time state feedback, reducing robustness. Another solution, Action Chunking [32], averages predictions over time for consistency but may yield suboptimal results by averaging dissimilar plans. Alternatively, [31] selects the closest trajectory to the previous one but incurs overhead from concurrently sampling multiple trajectories.
@@
Inference speedup can be naïvely achieved by minimizing diffusion steps at the cost of diversity, though techniques such as one-step diffusion [33] could mitigate these tradeoffs. Hybrid methods, which pair fast architectures for initial guesses with slower fine-tuning of models [34], balance speed and accuracy. Rolling or streaming strategies apply noise based on temporal proximity—less for immediate actions and more for distant ones—enabling faster generation of infinite sequences [9, 35]. Building upon this, we find that rolling addresses both motion consistency and speedup by maintaining diversity for future decisions while solidifying immediate ones.
III Method
@@
Figure 1: Framework of Diffuse-CLoC. The rolling scheme (Left) is implemented as a FIFO buffer, where at every timestep $t$, a state and action pair of pure noise is pushed into the denoiser, while the earliest clean pair is popped out and used to step the simulator. The denoiser architecture (Middle) denoises the buffer of states and actions with an increasing noise level along the trajectory. The observation $O_{t}$ is directly inpainted into the sequence, and only noisy future predictions are denoised. Notice that noisy states and actions have different shades, meaning that their noise levels $\bm{k_{s}}$, $\bm{k_{a}}$ can be different. The attention mask (Right) shows that state attends to all past and future states but masks out all actions, while action is causal attending to previous states and actions.
@@
Our approach to modeling complex motion behaviors builds upon denoising diffusion probabilistic models (DDPMs) [36]. At each timestep $t$, our model predicts a trajectory $\bm{\tau}_{t}=[\bm{a}_{t},\bm{s}_{t+1},\bm{a}_{t+1},\dots,\bm{s}_{t+H},\bm{a}_{t+H}]$ that contains the current action as well as the $H$ steps of state-action pairs in the future. Moreover, we let $\bm{O}_{t}$ be the observation history of state-action pairs of length $N$, as well as the current state, $\bm{s}_{t}$, i.e., ${\bm{O}_{t}=[\bm{s}_{t-N},\bm{a}_{t-N},\cdots,\bm{s}_{t}]}$.
We train a prediction network, $\hat{\bm{\tau}}_{t}$ = $x_{0,\theta}(\bm{\tau}_{t}^{\bm{k}},\bm{O}_{t},\bm{k})$ that aims to produce a clean trajectory, where $\bm{\tau}_{t}^{\bm{k}}$ is $\bm{\tau}$ with added Gaussian noise based on the noise level $\bm{k}$. We assign independent noise levels per timestep for states and actions along the trajectory, i.e. $\bm{k}=(\bm{k_{s}},\bm{k_{a}})$, where $\bm{k_{s}},\bm{k_{a}}\in\mathbb{R}^{N+H+1}$. We then follow Stochastic Langevin Dynamics [37] to sample iteratively,
$\bm{\tau}_{t}^{\bm{k}-1}=\alpha_{\bm{k}}(\bm{\tau}_{t}^{\bm{k}}-\gamma_{\bm{k}}\epsilon_{\theta}(\bm{\tau}_{t}^{\bm{k}},\bm{O}_{t},\bm{k})+\mathcal{N}(0,{\sigma_{\bm{k}}}^{2}\bm{I})),$ (1)
@@
$\mathcal{L}=MSE(x_{0,\theta}(\bm{\tau}_{t}^{\bm{k}},\bm{O}_{t},\bm{k}),\bm{\tau}_{t})$ (2)
Diffusion models, though trained unconditionally, can be adapted for conditional generation during inference. By Bayes’ rule, the denoiser’s learned score function $\nabla_{\bm{\tau}}\log p(\bm{\tau})$ [38] can be extended to the conditional form $\nabla_{\bm{\tau}}\log p(\bm{\tau}\mid c)=\nabla_{\bm{\tau}}\log p(\bm{\tau})+\nabla_{\bm{\tau}}\log p(c\mid\bm{\tau})$. Given a cost function $G_{\bm{\tau}}^{c}(\bm{\tau})$, a conditional probability $p(c\mid\bm{\tau})\propto\exp(-G_{\bm{\tau}}^{c}(\bm{\tau}))$ yields the posterior gradient $\nabla_{\bm{\tau}}\log p(c\mid\bm{\tau})=-\nabla_{\bm{\tau}}G_{\bm{\tau}}^{c}(\bm{\tau})$, which follows the cost’s descending direction. This approach, formally classifier guidance, conditions the diffuser using any cost function $G_{\bm{\tau}}^{c}(\bm{\tau})$ with a computable gradient $\nabla_{\bm{\tau}}G_{\bm{\tau}}^{c}(\bm{\tau})$.
@@
III-A Co-diffusing States and Actions

### v3

@@
Abstract.
We present Diffuse-CLoC, a guided diffusion framework for physics-based look-ahead control that enables intuitive, steerable, and physically realistic motion generation. While existing kinematics motion generation with diffusion models offer intuitive steering capabilities with inference-time conditioning, they often fail to produce physically viable motions. In contrast, recent diffusion-based control policies have shown promise in generating physically realizable motion sequences, but the lack of kinematics prediction limits their steerability. Diffuse-CLoC addresses these challenges through a key insight: modeling the joint distribution of states and actions within a single diffusion model makes action generation steerable by conditioning it on the predicted states. This approach allows us to leverage established conditioning techniques from kinematic motion generation while producing physically realistic motions. As a result, we achieve planning capabilities without the need for a high-level planner. Our method handles a diverse set of unseen long-horizon downstream tasks through a single pre-trained model, including static and dynamic obstacle avoidance, motion in-betweening, and task-space control. Experimental results show that our method significantly outperforms the traditional hierarchical framework of high-level motion diffusion and low-level tracking.
Character Animation, Unsupervised Reinforcement Learning, Diffusion Policy
submissionid: 1047journal: TOGjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 4publicationmonth: 8doi: 10.1145/3731206journalyear: 2025journalvolume: 44journalnumber: 4publicationmonth: 8doi: 10.1145/3731206ccs: Computing methodologies Procedural animationccs: Computing methodologies Motion path planning
Figure 1. Diffuse-CLoC, demonstrates impressive versatility across a wide range of unseen physics-based tasks using the same pre-trained diffusion model without the need for retraining or finetuning. These tasks include static and dynamic obstacle navigation with 16 characters, agile motion obstacle avoidance, gamepad control guided by a classifier, and physics-based motion inbetweening through inpainting.
1. Introduction
@@
Recent advances in kinematic motion diffusion models have shown promise in motion synthesis through conditioning on partial states (Shi et al., 2024), text (Tevet et al., 2023), and music (Tseng et al., 2023). Furthermore, the same kinematics motion diffusion model can be reused on unseen downstream tasks via inference-time conditioning such as classifier-guidance or inpainting (Karunratanakul et al., 2023; Cohan et al., 2024). Yet, these approaches fall short of achieving physical realism. To improve physical consistency, a tracking policy can be used to follow the kinematics motions, but ensuring robustness to domain shifts introduced by the kinematics generation remains a challenge. On the other hand, prior works (Chi et al., 2023; Truong et al., 2024; Huang et al., 2024a) have focused on predicting actions, ensuring physical realism without requiring extra components. However, unlike kinematic-based diffusion models, these approaches cannot adapt to unseen downstream tasks without retraining, as inference-time conditioning techniques are infeasible due to the lack of direct comparability between goals defined in state space and synthesized action trajectories.
@@
Our primary insight lies in modeling the joint distribution of states and actions within a single diffusion model, which enables action generation to be conditioned on the predicted states. Building on this insight, we develop a transformer-based diffusion architecture and an attention mechanism that includes non-causal attention for states and causal attention for actions. We further introduce a rolling scheme inspired by (Zhang et al., 2024) to improve consistency and speed in auto-regressive policy execution. Ultimately, Diffuse-CLoC enables the application of established techniques from kinematic motion generation, such as classifier guidance, to perform multiple unseen downstream tasks with physics-based characters, including waypoint navigation, dynamic obstacle avoidance such as other characters, and dynamic maneuvers like jumping over obstacles and crawling underneath them, as well as inpainting for motion in-betweening task (Fig. 1). Our primary contributions include:
@@ 2. Related Work
Figure 2. Three Formulations in Physics-based Control using Diffusion. (a) The factored distribution approach separately learns planning $p(s)$ and control $p(a|s)$, using kinematics planners (Karunratanakul et al., 2023; Tevet et al., 2023; Zhang et al., 2024) and tracking policies (Luo et al., 2023). Examples of this category are (Xie et al., 2023; Tevet et al., 2024; Ajay et al., 2023). (b) The model-free approach learns $p(a|s)$ only and does not allow inference-time planning (Truong et al., 2024; Chi et al., 2023; Huang et al., 2024a). (c) The joint distribution approach models $p(s,a)$ directly, enabling state-guided action generation. This includes (Janner et al., 2022) and Diffuse-CLoC, where Diffuse-CLoC demonstrates its effectiveness in complex character control tasks.
2.1. Physics-Based Character Animation
@@
Early physics-based character animation focused on tracking single motion trajectories (Peng et al., 2018). Subsequent work expanded this to capture motion distributions, such as AMP (Peng et al., 2021), though these approaches faced challenges in training stability and were limited to small datasets. As an alternative, latent skill embeddings were introduced, leveraging a VAE trained on large datasets (Yao et al., 2022, 2024; Won et al., 2022; Zhu et al., 2023), which can also be paired with a high-level controller to complete downstream tasks (Peng et al., 2022; Zhu et al., 2023; Won et al., 2022).
@@
Recent motion tracking controllers (Luo et al., 2023; Serifi et al., 2024b) have achieved notable success, leading to approaches that decompose the problem into two parts: a kinematic motion synthesis model and a universal tracking controller for the synthesized motion (Xie et al., 2023; Tevet et al., 2024). However, the tracking controller’s performance heavily depends on the quality of the kinematic trajectory, which can include artifacts like floating bodies, foot sliding, and penetration.
@@
To address this challenge, RobotMDM (Serifi et al., 2024a) fine-tunes a motion diffusion model using the value function of an RL tracker to produce more trackable motions. Similarly, (Han et al., 2024; Li et al., 2024a) refine diffusion models directly with PPO-based reinforcement learning guided by an RL tracker’s reward. Other methods instead fine-tune the tracking controller to handle artifacts created by the diffusion model (Tevet et al., 2024).
@@
Despite these advancements, trackers are often sensitive to out of distribution motions and require task-specific fine-tuning for applications like object interaction (Tevet et al., 2024). Furthermore, we demonstrate that even minor classifier guidance on kinematic trajectories can hurt the robustness of general tracking controllers.
2.2. Diffusion in Physics-based Control
@@
Diffusion policies have been successfully applied in character animation (Truong et al., 2024), manipulation (Chi et al., 2023), and locomotion (Huang et al., 2024a; Mothish et al., 2024) domains. However, its action-only generation makes inference-time conditioning impractical. In contrast, Diffuser (Janner et al., 2022) diffuses both states and actions, enabling classifier guidance through a reward formulation though exhibits limited robustness (Chen et al., 2024). Decision Diffuser (Ajay et al., 2023) improves robustness by similarly diffusing both states and actions, but discards actions during inference, relying instead on a trained inverse dynamics model. In contrast, our method enables both intuitive guidance directly in state space and high-quality conditional generation in action space.
2.3. Steering Motion Synthesis
@@
Classifier-free guidance enhances sample quality and control by interpolating between an unconditional and a conditional model during the generation process, without requiring an explicit classifier (Ho and Salimans, 2021). In character animation, this method has been used to synthesize motions with text conditioning (Truong et al., 2024; Tevet et al., 2023). However, the conditioning signals must be specified during training, and the model must be re-trained for any new conditioning signals.
@@
Classifier guidance incorporates conditional information after the training phase by utilizing the gradient of a pre-trained classifier (Dhariwal and Nichol, 2024) or cost function (Carvalho et al., 2023). This technique guides the diffusion model’s generative process, steering outputs toward the desired class or minimizing the loss. In character animation, it allows a single pre-trained kinematics model to address multiple downstream tasks, such as obstacle avoidance and reaching (Karunratanakul et al., 2023; Janner et al., 2022). Despite its success, classifier guidance has not yet been explored for dynamic tasks in underactuated robotic systems, such as physics-based character animation.
@@
Inpainting involves generating or reconstructing missing regions of the original signal by conditioning the diffusion process on the surrounding known context. In character animation, this technique is useful for generating trajectories with partially known states, such as root trajectories (Karunratanakul et al., 2023) or a set of desired joint trajectories (Tseng et al., 2023).
@@
Reinforcement learning has been employed as a high level planner to direct the diffusion process for downstream tasks (Shi et al., 2024). When future trajectory predictions are absent, a high-level planner is essential to strategize (Shi et al., 2024); however, our method can plan and foresee future events, allowing the accomplishment of downstream tasks without the need for a high-level planner.
2.4. Inference Consistency and Speed
@@
A key challenge for generative planners is ensuring that the generated plans do not conflict during rollouts. Frequent replanning with multimodal trajectory distributions can lead to alternating behavior modes, negatively affecting task performance (Römer et al., 2024). One approach is to execute a fixed action sequence before replanning (Chi et al., 2023), but this lacks real-time state feedback, reducing robustness. Another solution, Action Chunking (Zhao et al., 2023), averages predictions over time for consistency but may yield suboptimal results by averaging dissimilar plans. Alternatively, (Römer et al., 2024) selects the closest trajectory to the previous one but incurs overhead from concurrently sampling multiple trajectories.
@@
Inference speedup can be naïvely achieved by minimizing diffusion steps at the cost of diversity, though techniques such as one-step diffusion (Wang et al., 2024) could mitigate these tradeoffs. Hybrid methods, which pair fast architectures for initial guesses with slower fine-tuning of models (Li et al., 2024b), balance speed and accuracy. Rolling or streaming strategies apply noise based on temporal proximity—less for immediate actions and more for distant ones—enabling faster generation of infinite sequences (Zhang et al., 2024; Høeg et al., 2024). Building upon this, we find that rolling addresses both motion consistency and speedup by maintaining diversity for future decisions while solidifying immediate ones.
3. Method
@@
Figure 3. Framework of Diffuse-CLoC. The rolling scheme (Left) is implemented as a FIFO buffer, where at every timestep $t$, a state and action pair of pure noise is pushed into the denoiser, while the earliest clean pair is popped out and used to step the simulator. The denoiser architecture (Middle) denoises the buffer of states and actions with an increasing noise level along the trajectory. The observation $O_{t}$ is directly inpainted into the sequence, and only noisy future predictions are denoised. Notice that noisy states and actions have different shades, meaning that their noise levels $\bm{k_{s}}$, $\bm{k_{a}}$ can be different. The attention mask (Right) shows that state attends to all past and future states but masks out all actions, while action is causal attending to previous states and actions.
@@
Our approach to modeling complex motion behaviors builds upon denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020). At each timestep $t$, our model predicts a trajectory $\bm{\tau}_{t}=[\bm{a}_{t},\bm{s}_{t+1},\bm{a}_{t+1},\dots,\\ \bm{s}_{t+H},\bm{a}_{t+H}]$ that contains the current action as well as the $H$ steps of state-action pairs in the future. Moreover, we let $\bm{O}_{t}$ be the observation history of state-action pairs of length $N$, as well as the current state, $\bm{s}_{t}$, i.e., $\bm{O}_{t}=[\bm{s}_{t-N},\bm{a}_{t-N},\cdots,\bm{s}_{t}]$}.
We train a prediction network, $\hat{\bm{\tau}}_{t}$ = $x_{0,\theta}(\bm{\tau}_{t}^{\bm{k}},\bm{O}_{t},\bm{k})$ that aims to produce a clean trajectory, where $\bm{\tau}_{t}^{\bm{k}}$ is $\bm{\tau}$ with added Gaussian noise based on the noise level $\bm{k}$. We assign independent noise levels per timestep for states and actions along the trajectory, i.e. $\bm{k}=(\bm{k_{s}},\bm{k_{a}})$, where $\bm{k_{s}},\bm{k_{a}}\in\mathbb{R}^{N+H+1}$. We then follow Stochastic Langevin Dynamics (Welling and Teh, 2011) to sample iteratively,
(1) $\bm{\tau}_{t}^{\bm{k}-1}=\alpha_{\bm{k}}(\bm{\tau}_{t}^{\bm{k}}-\gamma_{\bm{k}}\epsilon_{\theta}(\bm{\tau}_{t}^{\bm{k}},\bm{O}_{t},\bm{k})+\mathcal{N}(0,{\sigma_{\bm{k}}}^{2}\bm{I})),$
@@
(2) $\mathcal{L}=MSE(x_{0,\theta}(\bm{\tau}_{t}^{\bm{k}},\bm{O}_{t},\bm{k}),\bm{\tau}_{t})$
Diffusion models, though trained unconditionally, can be adapted for conditional generation during inference. By Bayes’ rule, the denoiser’s learned score function $\nabla_{\bm{\tau}}\log p(\bm{\tau})$ (Song et al., 2021) can be extended to the conditional form $\nabla_{\bm{\tau}}\log p(\bm{\tau}\mid\bm{\tau}^{*})=\nabla_{\bm{\tau}}\log p(\bm{\tau})+\nabla_{\bm{\tau}}\log p(\bm{\tau}^{*}\mid\bm{\tau})$, where $\bm{\tau}^{*}$ is the optimal trajectory. Given a cost function $G_{\bm{\tau}}^{c}(\bm{\tau})$, we let the conditional probability of the optimal trajectory to be $p(\bm{\tau}^{*}\mid\bm{\tau})\propto\exp(-G_{\bm{\tau}}^{c}(\bm{\tau}))$. Then, the posterior gradient, $\nabla_{\bm{\tau}}\log p(\bm{\tau}^{*}\mid\bm{\tau})=-\nabla_{\bm{\tau}}G_{\bm{\tau}}^{c}(\bm{\tau})$, naturally aligns with the descending direction of the cost. This approach, formally classifier guidance, conditions the diffuser using any cost function $G_{\bm{\tau}}^{c}(\bm{\tau})$ with a computable gradient $\nabla_{\bm{\tau}}G_{\bm{\tau}}^{c}(\bm{\tau})$.
@@
3.1. Co-diffusing States and Actions